WRITE THE KCL EQUATIONS FOR THE TOP LEFT AND BOTTOM NODES. COMPARE YOUR FINDINGS

150 Ω 1.5 A ξ100 Ω ξ250 Ω
Answer #1

Guver cruits 1500 w 154 72 100.22 ² 250r How we need to find voltagevo, i andia using Keh . when we are applying kch we neeNow we should find the currents i, and is % = V1 - V2_ 12000 100 = Too F1 = 12A 12=vi-v2 180 + 250 - VI-V2 H00 = 129_0 400 =V: 120 V I: 300 mA i2 R2 X XMMT Multimeter-XMM1 120 V Probe21500 + Probe 1 11 V: 120 V R1 1000 vo(1)1.5A ~ : 1.20 A 11 R3 $25